In a circle whose diameter is 20 inches, A chord is 6 inches from the center. WHat is the length of the chord?: In a circle whose diameter is 20 inches, A chord is 6 inches from the center. WHat is the length of the chord?
Answer by ankor@dixie-net.com(4538)

In a circle whose diameter is 20 inches, A chord is 6 inches from the center. WHat is the length of the chord?
:
Draw this out and the method will apparent to you.
:
Note that inside the circle a triangle can be formed by the chord and 2 radii (10")
The triangle consists of two right triangles with sides 10", 6" and half the length of the chord, let it = a
:
Find "a" using pythag.
a^2 + 6^2 = 10^2
a^2 = 100 - 36
a^2 = 64
a = 8", therefore the length of the chord is 16"
:
:
did this make sense to you?